OpenStudy (anonymous):

Explain whether it is always true that a difference between two numbers is smaller than either of the numbers. Give an example to justify your answer.

OpenStudy (anonymous):

hey thechoco ur my fan but i never talk to u how come

OpenStudy (anonymous):

Let's say you want the difference of 12 and 6. 12-6 = 6 Not always 20 and 10? 20-10 = 10 10=10, so not always 4 and 3? 4-3 = 1 In this case, sometimes it is true.

OpenStudy (anonymous):

@cookies because you answered a question well

OpenStudy (anonymous):

oh thank you!! :)

OpenStudy (anonymous):

lol, no probs :)

OpenStudy (anonymous):

hmmmmm okay!! lol

OpenStudy (anonymous):

The chocoluver is right, but there's an even more dramatic example: 10 - (-20) = 30. In this example, the difference is actually *larger* than both numbers!